You find a zero coupon bond with a par value of $10,000 and 21 years to maturity. The yield to maturity on this bond is 4.3 percent. Assume semiannual compounding periods.

What is the price of the bond? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  Price $
Answer #1

Number of periods = 21 * 2 = 42

Rate = 4.3% / 2 = 2.15%

Price of bond = Face value / (1 + r)n

Price of bond = 10000 / (1 + 0.0215)42

Price of bond = 10000 / 2.4435

Price of bond = $4,092.50